# Debate or Vote: Which Yields Better Decisions in Multi-Agent Large Language Models?

### v1

@@ 1 Introduction
Throughout history, humans have relied on deliberation to resolve ambiguity, challenge assumptions, and seek better answers. From courtrooms and panels to scientific collaborations, group reasoning plays a central role in decision-making. This process—where individuals reflect, revise, and converge through interaction—has long been seen as a hallmark of intelligent behavior. Inspired by this, recent work has explored whether large language models (LLMs) might similarly benefit from structured interaction. Multi-Agent Debate (MAD) has emerged as a popular framework: multiple LLM agents are prompted to discuss a shared question, each updating their answer based on the responses of their peers bo2024reflective ; du2024improving ; chanchateval ; tang2024medagents ; wuautogen ; chenagentverse . The hope is that, like human deliberation, such interaction will improve reasoning and lead to better outcomes.
At its core, MAD integrates two key ingredients: the use of multiple agents (“Multi-Agent”) and their interaction through iterative discussions (“Debate”). Recent work has introduced increasingly sophisticated variants—ranging from diverse communication protocols chanchateval ; xiong2023examining ; liu2024groupdebate , designing efficient and effective system architectures bo2024reflective ; du2024improving ; liu2024dynamic ; li2024improving , and assigning varied roles or personas to agents chen2024reconcile ; liang2024encouraging ; wang2024unleashing . Despite these advances, the underlying mechanisms behind MAD’s effectiveness remain unclear. A natural step toward understanding MAD’s performance is to disentangle the contribution of each component—are the gains primarily due to meaningful communication between agents, or simply the result of aggregating multiple outputs? Answering this question is important because it informs whether the growing complexity of MAD design is justified by tangible benefits. If most of the performance gain stems from ensembling—i.e., aggregating diverse outputs from independent agents—then simpler methods like majority voting may suffice, avoiding additional computational and architectural overhead (see Figure 2 for visual comparison).
@@ 2 Preliminaries
is a collaborative framework in which multiple language model agents engage in structured interaction—typically in the form of iterative exchanges or discussions—to solve a task such as question answering or text generation bo2024reflective ; du2024improving ; chanchateval ; tang2024medagents ; wuautogen ; chenagentverse . In a typical MAD protocol, each agent independently generates an initial response and then engages in a series of debate rounds. At round $t$, an agent receives the original question along with responses from its peers at round t1t-1, prompting the model to update its response accordingly. This iterative process is designed to leverage diverse reasoning paths and peer wisdom, potentially enhancing the overall decision quality. After all rounds of debate, the final answer is typically derived through an aggregation mechanism, such as majority voting. Specific prompts are provided in Appendix B.1.
@@
where $\mathcal{N}(i)\subseteq\{1,\dots,N\}$ is the index set of neighbors, observable to agent $a_{i}$ (e.g., in the fully connected setting, $\mathcal{N}(i)=\{1,\dots,N\}$). The response update is given by:
@@
We adopt the simultaneous-talk protocol chanchateval , where all agents update in parallel based on the previous round’s responses. Following the common setup in prior works, we focus on homogeneous agent settings, i.e., all agents share the same underlying model architecture or behavior. This allows us to isolate the effect of inter-agent communication and contrast MAD directly with simple majority voting. Our goal is to contrast the performance of MAD against simple majority voting and assess whether iterative inter-agent communication provides measurable improvements beyond ensembling alone. We will extend to the heterogeneous setting in Section 6.
@@ 3.1 Experimental Setup
The key distinction among multi-agent debate methods typically lies in the design of the debate function $\mathcal{D}$, particularly in how agents communicate and the roles they assume. To comprehensively evaluate these variations, we consider the following representative approaches: (1) Decentralized MAD du2024improving , where each agent observes all other agents’ responses from the previous round. (2) Sparse MAD li2024improving , a variant of Decentralized MAD with a sparse communication topology to enhance efficiency. (3) Centralized MAD guo2024large , where a central agent aggregates peer responses and generates the updated response at each round. (4) Majority Voting, which selects the final answer by aggregating initial responses from multiple agents without any debate. This can be viewed as a special case with $T=0$. For all the multi-agent approaches, we adopt $N=5$ in our main comparison and will ablate on the effect of $N$. For single-agent baselines, we average across 5 independent runs.
@@
Following previous MAD literature, we focus on solving six natural language question answering tasks by conducting extensive evaluations across benchmark datasets: (1) Arithmetics, (2) Mathematical Reasoning (Grade School Math 8k cobbe2021gsm8k ), (3) Factual Question Answering (MMLU Professional Medicine and Formal Logics hendryckstest2021 ; hendrycks2021ethics ), (4) Natural Langauge Inference (HellaSwag zellers2019hellaswag ), (5) Commonsense Reasoning (CommonsenseQA talmor2019commonsenseqa ), and (6) Alignment Labeling (HH-RLHF bai2022training ), where we adopt the “AI labeler alignment” practice lee2024rlaif , similar to li2024improving . For fairness in comparison, all baselines are evaluated on the same data subsets. More details are provided in Appendix A.2.
@@ 3.2 Key Observations
In Table 1, we compare the performance of single-agent, MAD, and majority voting approaches across seven benchmark datasets using the Qwen2.5-7B-Instruct yang2024qwen2 and Llama3.1-8B-Instruct grattafiori2024llama models. Consistent with the typical choice of existing literature, we compare 2- and 3-round, along with a prolonged 5-round debate setting among five agents. Interestingly, while MAD consistently outperforms the single-agent baseline, it does not reliably surpass the much simpler majority voting strategy. Notably, in most cases, majority voting performs on par with MAD. To further assess the impact of model capacity, we additionally evaluate the more capable Qwen2.5-32B-Instruct model in Section 6. Although overall performance improves in the MAD setting, the majority voting strategy continues to account for most of the performance gains. These findings suggest that the effectiveness of MAD is largely driven by model ensembling, rather than the iterative debate process itself.
@@ 4 Theoretical Analysis
This result highlights the magnifying effect of majority voting: even when the correct answer is only marginally more likely than the alternative answers, the lower bound on the success probability asymptotically approaches 1 as $N$ scales. Notably, this holds even when $\bar{\theta}_{1}\ll\frac{1}{2}$, as long as it remains the most probable choice option. In practice, we recognize that MAD systems often operate with a small number of agents due to computational constraints. We provide a sharper analysis specialized to this realistic regime that applies to arbitrary $N$ in Theorem 1.A (Appendix D), without constraining its size. This complementary result provides insights into the reliability of majority voting in more practical, resource-constrained settings.
@@
As a result, we can prove that the evolution of agent beliefs forms a martingale process, with full proof in Appendix C.
@@
For agent $i$, let $p_{t}:=\bar{{\theta}}_{i,t}^{(1)}$ denote its belief in the correct answer at debate round $t$. Under Bayesian conjugacy,
$p_{t}:=\bar{\theta}_{i,t}^{(1)}=\frac{\alpha_{i,t}^{(1)}}{\sum_{j=1}^{K}\alpha_{i,t}^{(j)}}=\frac{\alpha_{i,t-1}^{(1)}+c_{i,t}^{(1)}}{\sum_{j=1}^{K}(\alpha_{i,t-1}^{(j)}+c_{i,t}^{(j)})}.$
Then, sequence $\{p_{t}\}_{t\geq 0}$ forms a martingale. That is, the expected belief at the next round equals the current belief:
$\mathbb{E}[p_{t}\mid p_{t-1},\ldots,p_{0}]=p_{t-1},\;\forall_{t\geq 0}.$
@@
This theorem highlights a fundamental property of multi-agent debate under Bayesian update of the DCM model: the agent’s belief in the correct answer behaves like a martingale—that is, its expected value remains unchanged across rounds. This result is closely related to the classical Pólya Urn scheme pemantle2007survey . Intuitively, this means that debate itself does not systematically improve or degrade an agent’s belief on average; instead, belief updates are driven entirely by the stochastic influence of peer responses. While some debate trajectories may lead to stronger belief in the correct answer (i.e., correction), others may lead to weaker belief (i.e., subversion). While these local fluctuations affect the posterior counts, the expected belief in the correct answer remains equal to the initial $p_{0}=\bar{\theta}_{i,0}^{(1)}$, without any debate. This implies that, under our theoretical model, debate alone does not necessarily improve the initial accuracy–majority voting accounts for the primary performance gains. Our theory thus aligns well with our empirical findings in Section 3.
@@
We empirically examine whether the sequence $\{p_{t}\}_{t\geq 0}$ exhibits martingale behavior. For each benchmark and debate round $t$, we estimate $p_{t}$ as the mean accuracy of the five agents [^1] Note that this quantity is different from the accuracy in Table 1, which is the ratio of correct responses after voting.. As shown in Figure 4, the resulting trajectories are essentially flat, which is consistent with the theoretical property that the expected value of a martingale remains unchanged over time. Raw values for mean accuracy are provided in Table 6 (Appendix E).
@@ 5.1 Belief Update by Biasing Towards Correct Signal
Table 2: Improved design of MAD. Model is Qwen2.5-7B-Instruct.
@@ 6 Extended Experiments to General Settings
To assess the generality of our findings, we extend our evaluation to more capable language models. Specifically, we test our setup on the Qwen2.5-32B-Instruct model yang2024qwen2 , using two representative tasks: GSM8K and HellaSwag. As shown in Table 4, the results confirm our earlier observations that the performance of Majority Voting remains comparable to that of multi-agent methods. This suggests that our claim is not limited to smaller models, but also holds in high-capacity LLMs.
@@
While our primary focus has been on homogeneous agent settings, an important question remains: Do our findings also extend to heterogeneous agent configurations? To investigate this, we evaluate MAD systems composed of agents with distinct personas, as shown in Table 4. Following the optimal persona sets identified for “college mathematics” and “clinical knowledge” via the “agent selection algorithm” introduced in liu2024dynamic , we construct diverse agent roles for each task. For GSM8K, the team includes a general-purpose “Assistant” alongside specialized roles: “Mathematician”, “Lawyer”, “Economist”, and “Programmer”. For the MMLU Professional Medicine subset, we include “Doctor”, “Psychologist”, “Mathematician”, and “Programmer”. In practice, we implement this by assigning each agent a system prompt that encodes a specific role or persona—e.g., Mathematician, Programmer, Lawyer—using the same prompt templates provided in liu2024dynamic  (see Appendix B.3 for the prompts). Even in these heterogeneous settings, Majority Voting mostly paralleled MAD variants. However, several MAD results on Pro. Med. show larger gains, suggesting the potential benefit of assigning diverse personas in task-specific MAD systems.
@@
In our previous experiments, we mainly focused on closed-ended question answering tasks, which are the primary focus of previous works on MAD. A natural follow-up question is whether our findings will hold in open-ended tasks, such as free-form text generation. To explore this, we evaluate MAD on a text summarization task using a subset of CNN/DailyMail dataset see2017get . Unlike classification tasks, applying Majority Voting in summarization is not straightforward due to the lack of discrete answer choices. Instead, we report the best-performing agent at each debate round, as shown in Table 5. Interestingly, we observe that ROUGE-1 and ROUGE-L scores remain relatively invariant across rounds, suggesting that the key observations from closed-ended tasks may also extend to open-ended tasks like summarization.
@@ 7 Related Works
Recently, there has been growing interest in multi-agent systems (MAS). Several survey papers have reviewed state-of-the-art LLM-based MAS approaches guo2024large ; tran2025multi ; yan2025beyond ; li2024survey . Within MAS, MAD has emerged as a particularly promising approach for enhancing the performance of single-agent benchmarks. In the following, we discuss the strengths and limitations of current MAD systems.
@@
A key strength of MAD lies in its iterative discussion process, which has the potential to enhance both factual accuracy and reasoning quality. Building on this paradigm, several works have proposed MAD-based approaches for a variety of tasks bo2024reflective ; du2024improving ; chanchateval ; tang2024medagents ; wuautogen ; chenagentverse . To further advance MAD systems, xiong2023examining introduced enhancements grounded in debate theory, while liprd developed the Peer Rank and Peer Discussion mechanisms to select appropriate agent pairs for debate. Many studies have focused on designing effective communication architectures and protocols to improve efficiency and effectiveness chanchateval ; liu2024groupdebate ; liu2024dynamic ; li2024improving ; phamlet ; zhang2024cut . Other works have emphasized the importance of diversity in MAD systems, leveraging heterogeneous LLM agents chen2024reconcile , injecting distinct personas into each agent liu2024dynamic ; liang2024encouraging ; wang2024unleashing , or enabling text generation with controlled diversity liubreaking ; chu2024exploring . Additionally, learning-based methods have been explored to optimize MAD dynamics liu2024dynamic ; estornellacc ; chen2024optima .
@@
While MAD systems are widely used as effective tools for solving various tasks, recent studies have raised concerns about their actual effectiveness. For instance, cemri2025multi conducted an in-depth analysis identifying 14 distinct failure modes in MAD systems, and zhang2025if found that MAD does not consistently outperform single-agent approaches. Similarly, huanglarge showed that LLM agents are not self-corrective enough for MAD to be successful, and smit2024should reported that MAD performs no better than advanced single-agent reasoning methods and highlighted its sensitivity to hyperparameters. wang2024rethinking echoed these concerns, showing that well-prompted single agents can sometimes outperform MAD. More specifically, xiong2023examining and zhang2025if observed occurrences of subverted or incorrect answers in MAD debates, while estornell2024multi showed that MAD systems often converge to the majority opinion, even when that opinion reflects common misconceptions. On a slightly different gear, benedikt2025voting compared various decision protocols and showed that multiple rounds of MAD actually decreases performance. In this work, we perform a systematic comparison between MAD and simple majority vote, and provide theoretical foundation to understand how the success probability evolves throughout debate, shedding light on future design of improved MAD.

### v2

@@ 1 Introduction
Throughout history, humans have relied on deliberation to resolve ambiguity, challenge assumptions, and seek better answers. From courtrooms and panels to scientific collaborations, group reasoning plays a central role in decision-making. This process—where individuals reflect, revise, and converge through interaction—has long been seen as a hallmark of intelligent behavior. Inspired by this, recent work has explored whether large language models (LLMs) might similarly benefit from structured interaction. Multi-Agent Debate (MAD) has emerged as a popular framework: multiple LLM agents are prompted to discuss a shared question, each updating their answer based on the responses of their peers [1, 2, 3, 4, 5, 6]. The hope is that, like human deliberation, such interaction will improve reasoning and lead to better outcomes.
At its core, MAD integrates two key ingredients: the use of multiple agents (“Multi-Agent”) and their interaction through iterative discussions (“Debate”). Recent work has introduced increasingly sophisticated variants—ranging from diverse communication protocols [3, 7, 8], designing efficient and effective system architectures [1, 2, 9, 10], and assigning varied roles or personas to agents [11, 12, 13]. Despite these advances, the underlying mechanisms behind MAD’s effectiveness remain unclear. A natural step toward understanding MAD’s performance is to disentangle the contribution of each component—are the gains primarily due to meaningful communication between agents, or simply the result of aggregating multiple outputs? Answering this question is important because it informs whether the growing complexity of MAD design is justified by tangible benefits. If most of the performance gain stems from ensembling—i.e., aggregating diverse outputs from independent agents—then simpler methods like majority voting may suffice, avoiding additional computational and architectural overhead (see Figure 2 for visual comparison).
@@ 2 Preliminaries
is a collaborative framework in which multiple language model agents engage in structured interaction—typically in the form of iterative exchanges or discussions—to solve a task such as question answering or text generation [1, 2, 3, 4, 5, 6]. In a typical MAD protocol, each agent independently generates an initial response and then engages in a series of debate rounds. At round $t$, an agent receives the original question along with responses from its peers at round t1t-1, prompting the model to update its response accordingly. This iterative process is designed to leverage diverse reasoning paths and peer wisdom, potentially enhancing the overall decision quality. After all rounds of debate, the final answer is typically derived through an aggregation mechanism, such as majority voting. Specific prompts are provided in Appendix B.1.
@@
where $\mathcal{N}(i)\subseteq\{1,\dots,N\}$ is the index set of neighbors observable to agent $a_{i}$, including itself (e.g., in the fully connected setting, $\mathcal{N}(i)=\{1,\dots,N\}$). The response update is given by:
@@
We adopt the simultaneous-talk protocol [3], where all agents update in parallel based on the previous round’s responses. Following the common setup in prior works, we focus on homogeneous agent settings, i.e., all agents share the same underlying model architecture or behavior. This allows us to isolate the effect of inter-agent communication and contrast MAD directly with simple majority voting. Our goal is to contrast the performance of MAD against simple majority voting and assess whether iterative inter-agent communication provides measurable improvements beyond ensembling alone. We will extend to the heterogeneous setting in Section 6.
@@ 3.1 Experimental Setup
The key distinction among multi-agent debate methods typically lies in the design of the debate function $\mathcal{D}$, particularly in how agents communicate and the roles they assume. To comprehensively evaluate these variations, we consider the following representative approaches: (1) Decentralized MAD [2], where each agent observes all other agents’ responses from the previous round. (2) Sparse MAD [10], a variant of Decentralized MAD with a sparse communication topology to enhance efficiency. (3) Centralized MAD [14], where a central agent aggregates peer responses and generates the updated response at each round. (4) Majority Voting, which selects the final answer by aggregating initial responses from multiple agents without any debate. This can be viewed as a special case with $T=0$. For all the multi-agent approaches, we adopt $N=5$ in our main comparison and will ablate on the effect of $N$. For single-agent baselines, we average across 5 independent runs.
@@
Following previous MAD literature, we focus on solving six natural language question answering tasks by conducting extensive evaluations across benchmark datasets: (1) Arithmetics, (2) Mathematical Reasoning (Grade School Math 8k [15]), (3) Factual Question Answering (MMLU Professional Medicine and Formal Logics [16, 17]), (4) Natural Langauge Inference (HellaSwag [18]), (5) Commonsense Reasoning (CommonsenseQA [19]), and (6) Alignment Labeling (HH-RLHF [20]), where we adopt the “AI labeler alignment” practice [21], similar to [10]. For fairness in comparison, all baselines are evaluated on the same data subsets. More details are provided in Appendix A.2.
@@ 3.2 Key Observations
In Table 1, we compare the performance of single-agent, MAD, and majority voting approaches across seven benchmark datasets using the Qwen2.5-7B-Instruct [22] and Llama3.1-8B-Instruct [23] models. Consistent with the typical choice of existing literature, we compare 2- and 3-round, along with a prolonged 5-round debate setting among five agents. Interestingly, while MAD consistently outperforms the single-agent baseline, it does not reliably surpass the much simpler majority voting strategy. Notably, in most cases, majority voting performs on par with MAD. To further assess the impact of model capacity, we additionally evaluate the more capable Qwen2.5-32B-Instruct model in Section 6. Although overall performance improves in the MAD setting, the majority voting strategy continues to account for most of the performance gains. These findings suggest that the effectiveness of MAD is largely driven by model ensembling, rather than the iterative debate process itself.
@@ 4 Theoretical Analysis
This result highlights the magnifying effect of majority voting: even when the correct answer is only marginally more likely than the alternative answers, the lower bound on the success probability asymptotically approaches 1 as $N$ scales. Notably, this holds even when $\bar{\theta}_{1}\ll\frac{1}{2}$, as long as it remains the most probable choice option. In practice, we recognize that MAD systems often operate with a small number of agents due to computational constraints. We provide a sharper analysis specialized to this realistic regime that applies to arbitrary $N$ in Theorem 1.A (Appendix C), without constraining its size. This complementary result provides insights into the reliability of majority voting in more practical, resource-constrained settings.
@@
Then, for an agent $i$ with neighbors $\mathcal{N}(i)$, let $p_{i,t}:=\bar{{\theta}}_{i,t}^{(1)}$ denote its belief in the correct answer at debate round $t$. Under Bayesian conjugacy,
$p_{i,t}:=\bar{\theta}_{i,t}^{(1)}=\frac{\alpha_{i,t}^{(1)}}{\sum_{j=1}^{K}\alpha_{i,t}^{(j)}}=\frac{\alpha_{i,t-1}^{(1)}+c_{i,t}^{(1)}}{\sum_{j=1}^{K}(\alpha_{i,t-1}^{(j)}+c_{i,t}^{(j)})}.$
This agent belief’s evolution throughout debate forms a martingale process (Theorem 2), with full proof in Appendix D.
@@
For any agent $i$ at round $t>0$, if
$\frac{1}{|\mathcal{N}(i)|}\sum_{j\in\mathcal{N}(i)}p_{j,t-1}=p_{i,t-1},$ (1)
then sequence $\{p_{i,t}\}_{t\geq 0}$ forms a martingale. That is, the expected belief at the next round equals the current belief:
$\mathbb{E}[p_{i,t}\mid\boldsymbol{\alpha}_{t-1}]=p_{i,t-1}.$
@@
This theorem highlights a fundamental property of multi-agent debate under Bayesian update of the DCM model: the agent’s belief in the correct answer behaves like a martingale—that is, its expected value remains unchanged across rounds. This result is closely related to the classical Pólya Urn scheme [24]. Intuitively, this means that debate itself does not systematically improve or degrade an agent’s belief on average; instead, belief updates are driven entirely by the stochastic influence of peer responses. While some debate trajectories may lead to stronger belief in the correct answer (i.e., correction), others may lead to weaker belief (i.e., subversion). While these local fluctuations affect the posterior counts, the expected belief in the correct answer remains equal to the initial $p_{i,0}=\bar{\theta}_{i,0}^{(1)}$, without any debate, when agents are homogeneous and fully-connected (Appendix D.3.1). This implies that, under our theoretical model, debate alone does not necessarily improve the initial accuracy–majority voting accounts for the primary performance gains. Our theory thus aligns well with our empirical findings in Section 3. Furthermore, in Appendix D.4, we extend our analysis to a more general setting in which each agent’s expected belief evolves over the course of the debate, such that the condition in Theorem 2 may no longer hold. In this context, we introduce the notion of “collective intelligence" and demonstrate that it forms a martingale process even with heterogeneous agents.
@@
We empirically examine whether the sequence $\{p_{i,t}\}_{t\geq 0}$ exhibits martingale behavior. For each benchmark and debate round $t$, we estimate $p_{t}$ as the mean accuracy of the five agents [^1] Note that this quantity is different from the accuracy in Table 1, which is the ratio of correct responses after voting.. As shown in Figure 4, the resulting trajectories are essentially flat, which is consistent with the theoretical property that the expected value of a martingale remains unchanged over time. Raw values for mean accuracy are provided in Table 6 (Appendix E).
@@ 5.1 Belief Update by Biasing Towards Correct Signal
Table 2: Improved design of MAD. The model is Qwen2.5-7B-Instruct.
@@ 6 Extended Experiments to General Settings
To assess the generality of our findings, we extend our evaluation to more capable language models. Specifically, we test our setup on the Qwen2.5-32B-Instruct model [22], using two representative tasks: GSM8K and HellaSwag. As shown in Table 4, the results confirm our earlier observations that the performance of Majority Voting remains comparable to that of multi-agent methods. This suggests that our claim is not limited to smaller models, but also holds in high-capacity LLMs.
@@
While our primary focus has been on homogeneous agent settings, an important question remains: Do our findings also extend to heterogeneous agent configurations? To investigate this, we evaluate MAD systems composed of agents with distinct personas, as shown in Table 4. Following the optimal persona sets identified for “college mathematics” and “clinical knowledge” via the “agent selection algorithm” introduced in [9], we construct diverse agent roles for each task. For GSM8K, the team includes a general-purpose “Assistant” alongside specialized roles: “Mathematician”, “Lawyer”, “Economist”, and “Programmer”. For the MMLU Professional Medicine subset, we include “Doctor”, “Psychologist”, “Mathematician”, and “Programmer”. In practice, we implement this by assigning each agent a system prompt that encodes a specific role or persona—e.g., Mathematician, Programmer, Lawyer—using the same prompt templates provided in [9] (see Appendix B.3 for the prompts). Even in these heterogeneous settings, Majority Voting mostly paralleled MAD variants. However, several MAD results on Pro. Med. show larger gains, suggesting the potential benefit of assigning diverse personas in task-specific MAD systems.
@@
In our previous experiments, we mainly focused on closed-ended question answering tasks, which are the primary focus of previous works on MAD. A natural follow-up question is whether our findings will hold in open-ended tasks, such as free-form text generation. To explore this, we evaluate MAD on a text summarization task using a subset of CNN/DailyMail dataset [25]. Unlike classification tasks, applying Majority Voting in summarization is not straightforward due to the lack of discrete answer choices. Instead, we report the best-performing agent at each debate round, as shown in Table 5. Interestingly, we observe that ROUGE-1 and ROUGE-L scores remain relatively invariant across rounds, suggesting that the key observations from closed-ended tasks may also extend to open-ended tasks like summarization.
@@ 7 Related Works
Recently, there has been growing interest in multi-agent systems (MAS). Several survey papers have reviewed state-of-the-art LLM-based MAS approaches [14, 26, 27, 28]. Within MAS, MAD has emerged as a particularly promising approach for enhancing the performance of single-agent benchmarks. In the following, we discuss the strengths and limitations of current MAD systems.
@@
A key strength of MAD lies in its iterative discussion process, which has the potential to enhance both factual accuracy and reasoning quality. Building on this paradigm, several works have proposed MAD-based approaches for a variety of tasks [1, 2, 3, 4, 5, 6]. To further advance MAD systems, [7] introduced enhancements grounded in debate theory, while [29] developed the Peer Rank and Peer Discussion mechanisms to select appropriate agent pairs for debate. Many studies have focused on designing effective communication architectures and protocols to improve efficiency and effectiveness [3, 8, 9, 10, 30, 31]. Other works have emphasized the importance of diversity in MAD systems, leveraging heterogeneous LLM agents [11], injecting distinct personas into each agent [9, 12, 13], or enabling text generation with controlled diversity [32, 33]. Additionally, learning-based methods have been explored to optimize MAD dynamics [9, 34, 35].
@@
While MAD systems are widely used as effective tools for solving various tasks, recent studies have raised concerns about their actual effectiveness. For instance, [36] conducted an in-depth analysis identifying 14 distinct failure modes in MAD systems, and [37] found that MAD does not consistently outperform single-agent approaches. Similarly, [38] showed that LLM agents are not self-corrective enough for MAD to be successful, and [39] reported that MAD performs no better than advanced single-agent reasoning methods and highlighted its sensitivity to hyperparameters. [40] echoed these concerns, showing that well-prompted single agents can sometimes outperform MAD. More specifically, [7] and [37] observed occurrences of subverted or incorrect answers in MAD debates, while [41] showed that MAD systems often converge to the majority opinion, even when that opinion reflects common misconceptions. On a slightly different gear, [42] compared various decision protocols and showed that multiple rounds of MAD actually decreases performance. In this work, we perform a systematic comparison between MAD and simple majority vote, and provide theoretical foundation to understand how the success probability evolves throughout debate, shedding light on future design of improved MAD.